# Circle packing and Riemann uniformization of random triangulations in an ergodic scale-free environment

### 06 Mar 2026 v1

@@ 4.1 Regularity estimates for the Riemann surface
We note that the proofs of Lemmas 4.3-4.7 can be adapted to the setting of circle packings. Therefore, by repeating the arguments in this subsection, it might be possible to give an alternative proof of Proposition 3.11 without using the vertex extremal length in [GJN19].
@@ 4.2 Existence of the harmonic corrector
Let $\mathcal{H}$ be a random cell configuration satisfying all the conditions in Theorem 1.8 and let $M(\mathcal{H})$ be the Riemann surface constructed from the map associated with $\mathcal{H}$. In this section we will define a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ as the limit of functions $\phi_{m}:M(\mathcal{H})\to\mathbb{C}$ for $m\in\mathbb{N}$, which are defined as certain harmonic extensions of an appropriately chosen function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$. This idea is similar to [GMS22, Section 2.3] and the key difference is that we work with (continuum) harmonic functions on $M(\mathcal{H})$ instead of discrete harmonic functions on the graph $\mathcal{H}$.
The function $\phi_{0}$ sends each vertex on $M(\mathcal{H})$ to a point in the associated cell of $\mathcal{H}$ and is otherwise piecewise linear. More precisely, the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$ is defined as follows. Given a realization of $\mathcal{H}$, for each cell $H\in\mathcal{H}$, we independently sample a point $\widetilde{c}(H)$ according to the probability measure proportional to the Lebsegue measure on $H$. For each triangular face $\Delta$ in $M(\mathcal{H})$ formed by the cells $H_{1},H_{2},H_{3}\in\mathcal{H}$, consider the triangle $\Delta^{\prime}$ in $\mathbb{C}$ formed by connecting $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ using straight line segments, and define $\phi_{0}|_{\Delta}:\Delta\to\Delta^{\prime}$ to be the unique linear function sending the vertices in $\Delta$ to their counterpart in $\Delta^{\prime}$. This defines a continuous and piecewise linear function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$. See Figure 8 for an illustration.
@@
We will also need the following topological properties on $\phi_{0}$.
Lemma 4.15.
We have $\phi_{0}(M(\mathcal{H}))=\mathbb{C}$.
Proof.
Pick $k$ large enough such that by Definition 1.4, the 4 subgraphs of $\mathcal{H}$ formed by the cells intersecting the 4 sides of $S_{k}$ are connected. Since a cell intersecting a vertex of $S_{k}$ is contained in (at least) two of these subgraphs, we get that the set of cells intersecting a side of $S_{k}$ forms a connected graph. By further applying Lemma 2.6, we can find a chain of cells $H_{1}\sim H_{2}\sim\cdots\sim H_{n}\sim H_{1}$ in this connected graph such that if we connect $\widetilde{c}(H_{j})$ and $\widetilde{c}(H_{j+1})$ using straight line segments for j=1,,nj=1,...,n with $n+1$ identified with 1, then this creates a loop $\gamma_{n}$ surrounding the square $\widetilde{S}_{k}$ concentric to $S_{k}$ with side length $|S_{k}|/2$, and the winding number of $\gamma_{n}$ around each point in $\widetilde{S}_{k}$ is equal to 1. If we consider the vertices for $H_{1},...,H_{n}$ on the surface $M(\mathcal{H})$, then one can find a loop $\widetilde{\gamma}_{n}$ on $M(\mathcal{H})$ connecting these vertices with straight line segments such that $\phi_{0}(\widetilde{\gamma}_{n})=\gamma_{n}$. Since $M(\mathcal{H})$ is simply connected, we can continuously shrink $\widetilde{\gamma}_{n}$ to a single point $x$. If there exists a point $w\in\widetilde{S}_{k}\backslash\phi_{0}(M(\mathcal{H}))$, then under $\phi_{0}$ image, all of these shrinking loops will have winding number 1 around $w$. On the other hand, for loops sufficiently close to $x$, its image under $\phi_{0}$ has winding number 0 around $w$ since $w\notin\phi_{0}(M(\mathcal{H}))$. This is a contradiction, which implies that almost surely for large enough $k$, $\widetilde{S}_{k}\subseteq\phi_{0}(\mathcal{H})$, which completes the proof. ∎
@@ 4.3 Sublinearity of the corrector
The aim of this subsection is to prove Proposition 4.20 below. Recall the notation $M(S)$ introduced at the beginning of Section 4.1. We start with the fact that the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ is small when $m$ is large (Lemma 4.22), and for fixed $m$, the maximal value of $|\phi_{m}(z)-\phi_{0}(z)|$ on $M(S_{k})$ has order $o_{k}(|S_{k}|)$ as $k\to\infty$ (Lemma 4.23). Then we transfer the bound on the the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ to a pointwise bound for $|\phi_{0}-\phi_{\infty}|$. The structure of the proof is similar as in [GMS22, Section 2.4], but we need to use a different setup when we conclude the proof of Proposition 4.20 where we we work with the uniformization map $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ as in Lemmas 4.3 and 4.4, and we also need to use regularity results from Section 4.1. We refer to the paragraph above the latter proof for a proof outline.
@@
Let $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ be the uniformization map defined as in Lemma 4.3 and Lemma $4.$ Using the conformal invariance of Dirichlet energy,
@@
We prove the decomposition of $\widetilde{\phi}_{\infty}$ as follows. First we prove that the harmonic function $\widetilde{\phi}_{\infty}$ must be a polynomial (Lemma 4.24). This follows from Proposition 4.20 as well as the polynomial growth property in Lemma 4.10. Then in Lemmas 4.254.26 and 4.27, we prove that this polynomial must have degree 1, by arguing that if this is not the case, then one can find cells $H_{1},H_{2}$ which are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.3 and Lemma 4.6. Next we use the ergodic theorem for cell configurations (Lemma 2.4) to prove that for $a,b\in\mathbb{R}$, the average Dirichlet energy of $a\,\mathrm{Re}\,\phi_{\infty}+b\,\mathrm{Im}\,\phi_{\infty}$ is deterministic (Lemma 4.29). This implies that for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$, if we view $\widetilde{\phi}_{\infty}$ as a $2\times 2$ matrix, then $\langle{\widetilde{\phi}_{\infty}^{\mathrm{T}}\mathbf{u}},{\widetilde{\phi}_{\infty}^{\mathrm{T}}\mathbf{v}}\rangle$ is deterministic up to constant (Lemma 4.30), and the desired decomposition is a consequence of standard linear algebra (Lemma 4.31). Finally in Lemma 4.33, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.6, then $\mathbf{A}$ must be a rotation.
@@
In order to conclude the proof we need to argue that the right side of the last indented equation is finite and that the left side does not change if the domain of integration is changed to be consistent with (4.77). Fix $\delta\in(0,1)$, and consider a square $S\subseteq S_{k}$ of side length at least $\delta|S_{k}|$. For each $z$ contained in some $H\in\mathcal{H}(S)$, if $z$ is inside the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ as in the definition of $\phi_{0}$, then for some $j$, $H_{j}\cap S\neq\emptyset$, and using Lemmas 2.6 and 4.1, it follows that a.s. for large enough $k$, $H_{1},H_{2},H_{3}\subseteq\mathring{S}^{1.5|S|}$ and thus $\phi_{0}^{-1}(z)\subseteq M(\mathring{S}^{2|S|})$. Therefore, a.s. for large $n$, for each square $S\subseteq S_{k}$ of side length at least $\delta|S|$,
@@
5 Extensions of Theorems 1.7 and 1.8
@@
In this section, we explain several extensions of Theorems 1.7 and 1.8. In Section 5.1, we prove Theorem 5.1, which is a version of Theorems 1.7 and 1.8 with a weaker notion of the line connectivity property in Definition 1.4. In Section 5.2, we prove an extension of Theorem 1.7 for a version of circle packing for general planar maps. In Section 5.3, we prove an extension of Theorem 1.8 for $p$-angulations with $p\in\{4,5,\dots\}$.
5.1 An extension with weaker notion of line connectivity
The following version of Theorem 1.7 and Theorem 1.8, which uses a weaker version of the line connectivity in Definition 1.4, will be used in our future work.
Theorem 5.1.
Suppose $\mathcal{H}$ is a cell configuration that satisfies all the conditions in Theorem 1.7 or Theorem 1.8 except for the line connectivity condition in Definition 1.4. Suppose there exists a cell configuration $\mathcal{H}^{\rm large}$ satisfying the condition in Definition 1.3 and the moment bound (1.3) along with the following properties:
  1. 1.
  2. 2.
  3. 3.
Then the conclusions in Theorem 1.7 or Theorem 1.8 continue to hold.
Theorem 5.1 can be viewed as a generalization of Theorem 1.7 and Theorem 1.8 since one can set $\mathcal{H}^{\rm large}=\mathcal{H}$ in the latter setting.
Proof.
Most of the proof is identical to that of Theorem 1.7 and Theorem 1.8, and we will only explain the differences. Note that the arguments in Section 2.2 do not require the line connectivity property in Definition 1.4 or the almost planarity condition in Defininition 1.5, so the conclusions there continue to hold. In particular, by Lemma 2.6 applied to $\mathcal{H}^{\rm large}$ and condition (ii) in the theorem statement, the almost planarity condition in Definition 1.5 holds for $\mathcal{H}^{\rm large}$ as well.
In the setting of circle packing, the arguments in Section 3 do not directly use the line connectivity property in Definition 1.4. The property is used implicitly as we carry the proof of [GMS22, Theorem 1.16]; in the proof in [GMS22], this property is used only in the proof of [GMS22, Lemma 2.19], in the paragraph right after [GMS22, Eq. (2.46)]. To substitute for that, Conditions (ii) and (iii) guarantee that a.s. for large enough $k$ and horizontal or vertical line segments $\ell\subseteq S_{k}$ of length $S_{k}$, for any function $g:\mathcal{H}^{\rm large}\to\mathbb{R}$ and $H_{1},H_{2}\in\mathcal{H}(\ell)$,
The corresponding quantity on the right side can be bounded exactly as [GMS22] since the arguments there only rely on the inputs from Section 2.2, which do not require the line connectivity property.
[…]
Then  (4.11) holds with $L^{\rm large}(\rho)$ instead of $L(\rho)$ using identical arguments applied to $\mathcal{H}^{\rm large}$, and (4.5) follows from identical arguments using conditions (ii) and (iii). For the proof of Lemma 4.15, using conditions (ii) and (iii) along with Lemma 2.6, one can find cells $H_{1}\sim\cdots H_{n}\sim H_{1}$ with identical properties and the conclusion still holds. The arguments for Lemma 4.26 are identical, except that at the final step we use that by Lemma 4.1 the set on the right hand side of (4.69) is contained within $\varphi(M(w+[-\delta 2^{k},\delta 2^{k}]^{2}))$, which is a connected subset of the set in (4.66). The rest of the proofs in Sections 4.1-4.4 do not require the line connectivity property and can be copied to the setting of Theorem 5.1. ∎
5.2 An Extension of Theorem 1.7 to general planar maps
@@ 5.1 An extension with weaker notion of line connectivity
Let $\mathcal{M}$ be an one-ended 2-connected simple planar map with face set $\mathcal{F}\mathcal{M}$. Recall from Section 1.4 that one can add a vertex $v_{f}$ to every face $f\in\mathcal{F}\mathcal{M}$, and connect the vertices $v_{f}$ to the vertices incident to $f$, to get a triangulation $\mathcal{T}_{\mathcal{M}}$. The triangulation $\mathcal{T}_{\mathcal{M}}$ will be simple since we required $\mathcal{M}$ to be 2-connected. We define the circle packing of $\mathcal{M}$ to be the circle packing for $\mathcal{T}_{\mathcal{M}}$ minus the disks for the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$. In this section we prove a variant of Theorem 1.7 in this setting.
Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ as above, and let $\mathcal{T}_{\mathcal{M}}$ be the triangulation generated from $\mathcal{M}$. We write $v_{H}\sim f$ if the vertex $v_{H}$ for $H$ in the associated map $\mathcal{M}$ is incident to the face $f\in\mathcal{F}\mathcal{M}$. Let
be the union of cells whose vertex is on the same face with $v_{H_{0}}$ and let $d_{0}$ be the maximal degree of the faces incident to $v_{H_{0}}$.
Theorem 5.2.
Let $\mathcal{H}$ be a cell configuration with the associated map $\mathcal{M}$ being an one-ended 2-connected simple planar map. Assume $\mathcal{H}$ satisfies all the conditions in Definitions 1.3, 1.4 and 1.5. Further assume
Then the conclusions in Theorem 1.7 holds for the circle packing of $\mathcal{M}$ defined as above.

### 06 May 2026 v2

@@ 4.1 Regularity estimates for the Riemann surface
We note that the proofs of Lemmas 4.2-4.6 can be adapted to the setting of circle packings. Therefore, by repeating the arguments in this subsection, it might be possible to give an alternative proof of Proposition 3.9 without using the vertex extremal length in [GJN19].
@@ 4.2 Existence of the harmonic corrector
Let $\mathcal{H}$ be a random cell configuration satisfying all the conditions in Theorem 1.7 and let $M(\mathcal{H})$ be the Riemann surface constructed from the map associated with $\mathcal{H}$. Recall the function $\phi_{0}$ defined in Section 2.2. In this section we will define a harmonic function $\phi_{\infty}:M(\mathcal{H})\to\mathbb{C}$ as the limit of functions $\phi_{m}:M(\mathcal{H})\to\mathbb{C}$ for $m\in\mathbb{N}$, which are defined as certain harmonic extensions of the function $\phi_{0}:M(\mathcal{H})\to\mathbb{C}$. This idea is similar to [GMS22, Section 2.3] and the key difference is that we work with (continuum) harmonic functions on $M(\mathcal{H})$ instead of discrete harmonic functions on the graph $\mathcal{H}$.
@@ 4.3 Sublinearity of the corrector
The aim of this subsection is to prove Proposition 4.18 below. Recall the notation $M(S)$ introduced at the beginning of Section 4.1. We start with the fact that the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ is small when $m$ is large (Lemma 4.20), and for fixed $m$, the maximal value of $|\phi_{m}(z)-\phi_{0}(z)|$ on $M(S_{k})$ has order $o_{k}(|S_{k}|)$ as $k\to\infty$ (Lemma 4.21). Then we transfer the bound on the the Dirichlet energy per unit area of $\phi_{m}-\phi_{\infty}$ to a pointwise bound for $|\phi_{0}-\phi_{\infty}|$. The structure of the proof is similar as in [GMS22, Section 2.4], but we need to use a different setup when we conclude the proof of Proposition 4.18 where we we work with the uniformization map $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ as in Lemmas 4.2 and 4.3, and we also need to use regularity results from Section 4.1. We refer to the paragraph above the latter proof for a proof outline.
@@
Let $\varphi_{k}:M(S_{k})\to|S_{k}|\mathbb{D}$ be the uniformization map defined as in Lemma 4.2 and Lemma 4.3. Using the conformal invariance of Dirichlet energy,
@@
We prove the decomposition of $\widetilde{\phi}_{\infty}$ as follows. First we prove that the harmonic function $\widetilde{\phi}_{\infty}$ must be a polynomial (Lemma 4.22). This follows from Proposition 4.18 as well as the polynomial growth property in Lemma 4.9. Then in Lemmas 4.234.24 and 4.25, we prove that this polynomial must have degree 1, by arguing that if this is not the case, then one can find cells $H_{1},H_{2}$ which are close in $\mathcal{H}$ and have macroscopic distance in $\varphi(M(\mathcal{H}))$, which contradicts with the continuity as in Lemma 4.2 and Lemma 4.5. Next we use the ergodic theorem for cell configurations (Lemma 2.4) to prove that for $a,b\in\mathbb{R}$, the average Dirichlet energy of $a\,\mathrm{Re}\,\phi_{\infty}+b\,\mathrm{Im}\,\phi_{\infty}$ is deterministic (Lemma 4.27). This implies that for any $\mathbf{u},\mathbf{v}\in\mathbb{R}^{2}$, if we view $\widetilde{\phi}_{\infty}$ as a $2\times 2$ matrix, then $\langle{\widetilde{\phi}_{\infty}^{\mathrm{T}}\mathbf{u}},{\widetilde{\phi}_{\infty}^{\mathrm{T}}\mathbf{v}}\rangle$ is deterministic up to constant (Lemma 4.28), and the desired decomposition is a consequence of standard linear algebra (Lemma 4.29). Finally in Lemma 4.31, we prove that if we further assume the rotation invariance in law of $\mathcal{H}$ as in Definition 1.5, then $\mathbf{A}$ must be a rotation.
@@
In order to conclude the proof we need to argue that the right side of the last indented equation is finite and that the left side does not change if the domain of integration is changed to be consistent with (4.77). Fix $\delta\in(0,1)$, and consider a square $S\subseteq S_{k}$ of side length at least $\delta|S_{k}|$. For each $z$ contained in some $H\in\mathcal{H}(S)$, if $z$ is inside the triangle formed by $\widetilde{c}(H_{1}),\widetilde{c}(H_{2}),\widetilde{c}(H_{3})$ as in the definition of $\phi_{0}$, then for some $j$, $H_{j}\cap S\neq\emptyset$, and using Lemmas 2.6 and 2.8, it follows that a.s. for large enough $k$, $H_{1},H_{2},H_{3}\subseteq\mathring{S}^{1.5|S|}$ and thus $\phi_{0}^{-1}(z)\subseteq M(\mathring{S}^{2|S|})$. Therefore, a.s. for large $n$, for each square $S\subseteq S_{k}$ of side length at least $\delta|S|$,
@@
5 Extensions to general planar maps
@@
In this section, we prove Theorem 1.8 and Theorem 1.9, extending Theorem 1.6 and Theorem 1.7 to the setting of general planar maps. In Section 5.3, we prove Theorem 5.5, which is a version of Theorems 1.6 and 1.7 with a weaker notion of the line connectivity property in Definition 1.4.
5.1 Proof of Theorem 1.8
@@
Let $\mathcal{M}$ be a 2-connected simple planar map with whole-plane topology and face set $\mathcal{F}\mathcal{M}$. Recall from Section 1.4 that one can add a vertex $v_{f}$ to every face $f\in\mathcal{F}\mathcal{M}$, and connect the vertices $v_{f}$ to the vertices incident to $f$, to get a triangulation $\mathcal{T}_{\mathcal{M}}$. The triangulation $\mathcal{T}_{\mathcal{M}}$ will be simple since we required $\mathcal{M}$ to be 2-connected. We define the circle packing of $\mathcal{M}$ to be the circle packing for $\mathcal{T}_{\mathcal{M}}$ minus the disks for the vertices $\{v_{f}:f\in\mathcal{F}\mathcal{M}\}$. In this section we prove a variant of Theorem 1.6 in this setting.